OpenStudy (anonymous):

What is th slope of the line that passes through (2,5) and (-1,5)?

OpenStudy (anonymous):

slope = y2 - y1 / x2 - x1

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (5-5)/(-1-2) m = 0/(-3) m = 0 So the slope of the line that passes through the two points is 0